OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 25?

OpenStudy (anonymous):

Since \[f^{-1}(f(x)) = x\]\[f(f^{-1}(f(x)))=f(x)\]Let \(u=f(x)\)\[ f(f^{-1}(u))=u \]

OpenStudy (anonymous):

\[ f(f^{-1}(u))=(f^{-1}(u))^2-25=u \]

OpenStudy (anonymous):

At this point you solve for \(f^{-1}(u)\): \[\begin{split} (f^{-1}(u))^2&=u+25\\ |f^{-1}(u)| &= \sqrt{u+25} \\ f^{-1}(u)&=\pm \sqrt{u+25} \end{split} \]

OpenStudy (anonymous):

Unfortunately, this isn't a function because it has two outputs, and functions only have on e output.